OpenStudy (anonymous):

Solve for x: 3 − (2x − 5) < −4(x + 2) x < −8 x > −8 x < −3 x > −3

OpenStudy (jacob902):

heres an example Solve - 3(2x-5)-4x=3(x-2)+x?

OpenStudy (jacob902):

-3(2x-5)-4x = 3(x-2)+x --- 1st multiply the brackets, so -6x+15-4x = 3x-6+x ---- now move the 'x' to the right no. to the left -6x-4x-3x-x = -6-15 (sym.change, so + turn -, & - into + -14x = -21 (just add them) x = -21/-14 x = 3/2 = 1.5

OpenStudy (anonymous):

@Jacob902 X = 1.5?

OpenStudy (anonymous):

@Jacob902 Is there a formula or like a way to do this that you can explain to me?

OpenStudy (laughoutloud):

3 − (2x − 5) < −4(x + 2)---Distribute out the Negative sign on the left side, and the -4 on the right side

OpenStudy (laughoutloud):

3 - 2x +5 < -4x - 8x

OpenStudy (laughoutloud):

sorry i meant 3 - 2x +5 < -4 - 8

OpenStudy (laughoutloud):

3 - 2x + 5 < -4x - 8 2x< -16 x<-8

OpenStudy (jacob902):

yep